[Brahms-l] Shift report 20050313 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Mar 13 2005 - 08:02:18 EST
Shifter leader was kh
Shifters were 


Had beam nearly all night.
Runs achieved:
Run   tr2     tr3
FS 3 deg 1/2A, MRS 90 700B
13757   61k   52k
13760   80k   82k
13761   60k   56k
13762   56k   49k
0 field FS runs
13763   24k    7k
13764   29k    9k
FS 6 deg 1/4A, MRS 90 700B
13766 in progress at shift end

Magnet crew changed D2 tap to tap4 (see elog for details on how 
this was chosen

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Mar 13 08:02:36 2005

This archive was generated by hypermail 2.1.8 : Sun Mar 13 2005 - 08:02:50 EST